Paunch Burger has a beta of 1.2 and just paid a dividend of $2.30 that is expected to grow at 3.2%. If the risk-free rate is 3% and the market risk premium is 6%, what should be the price of the stock?

A. $69.81

B. $39.69

C. $32.86

D. $33.91
Answer #1

Required rate=risk free rate+Beta*market risk premium

=3+(1.2*6)=10.2%

Current price=D1/(Required return-Growth rate)

=(2.3*1.032)/(0.102-0.032)

=$33.91(Approx).